OpenStudy (anonymous):

PLEASE FIND THE ANSWER: If the midpoint between (18, y) and (20, -15) is (19, -5), find the value of y.

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@LunyMoony

OpenStudy (anonymous):

please find the answer for me medal will be awarded I'm in a time crunch

OpenStudy (misty1212):

\[\frac{y-15}{2}=-5\] solver for \(y\)

OpenStudy (anonymous):

im not sure how to ddl that..

OpenStudy (anonymous):

@misty1212

OpenStudy (misty1212):

multiply both sides by 2, then add 15 to both sides are the two necessary steps

OpenStudy (anonymous):

so y=5?

OpenStudy (anonymous):

@misty1212

OpenStudy (misty1212):

yes dear that is the right answer

OpenStudy (anonymous):

Find the coordinates of the midpoint of if Jk (-9, 5) and K(21, -7).

OpenStudy (anonymous):

what about this? @misty1212

OpenStudy (anonymous):

thank u btw?:)

OpenStudy (misty1212):

you are welcome this is not too hard average , that is add and divide by two \[(\frac{-9+21}{2},\frac{5-7}{2})\]

OpenStudy (anonymous):

im not sure what you mean.. add them separately ?

OpenStudy (misty1212):

yeah the x's to the x's and the y's to the y's like i wrote above

OpenStudy (anonymous):

12/2 , -2/2 ...?

OpenStudy (misty1212):

lol yeah, but i think maybe you can do better right?

OpenStudy (anonymous):

idk.. I'm solo bad at this unit!!

OpenStudy (anonymous):

/unconfident with this I'm always second guessing myself

OpenStudy (misty1212):

that is ok you are doing fine

OpenStudy (anonymous):

but thAt would be the answer

OpenStudy (misty1212):

yeah but you need to write \\[(6,-1)\]

OpenStudy (anonymous):

ok would it be -6,1 because the 2 is negative?

OpenStudy (misty1212):

huh?

OpenStudy (anonymous):

for the answer

OpenStudy (misty1212):

what do you mean "cause the 2 is negative"? i mean the answer is NO, it is definitely \((6,-1)\)

OpenStudy (anonymous):

ok ur prob right lol thanks againn!

OpenStudy (misty1212):

you are welcome dear